OpenStudy (anonymous):

The cost of $500,000 worth of 20-year term life insurance for Derek is $96.21 per month. If Derek's employer covers 85% of this cost, how much is deducted from Derek's gross income per year for life insurance.

OpenStudy (anonymous):

(85/100) * 96.21=81.7785 in a year there are 12 months, so amount cut from Derek's gross income per year = 12 * 81.7785=$981.34

OpenStudy (anonymous):

fatma can you text me?